OpenStudy (anonymous):

the sqaure root of 243

OpenStudy (anonymous):

15.58 :)

OpenStudy (anonymous):

^incorrect rounding

OpenStudy (campbell_st):

what about \[\sqrt{243} = \sqrt{81 \times 3} = \sqrt{81} \times \sqrt{3}\] now just simplify from here...

OpenStudy (anonymous):

tru

OpenStudy (morganchavez):

15.28 rounded it would be 15.3

OpenStudy (morganchavez):

right ? @ButterflyHope

OpenStudy (butterflyhope):

Yep I'm pretty sure :)

OpenStudy (morganchavez):

kk thx